# Mutually-Aware Feature Learning for Few-Shot Object Counting

### v1

@@ Abstract
Few-shot object counting has garnered significant attention for its practicality as it aims to count target objects in a query image based on given exemplars without the need for additional training. However, there is a shortcoming in the prevailing extract-and-match approach: query and exemplar features lack interaction during feature extraction since they are extracted unaware of each other and later correlated based on similarity. This can lead to insufficient target awareness of the extracted features, resulting in target confusion in precisely identifying the actual target when multiple class objects coexist. To address this limitation, we propose a novel framework, Mutually-Aware FEAture learning (MAFEA), which encodes query and exemplar features mutually aware of each other from the outset. By encouraging interaction between query and exemplar features throughout the entire pipeline, we can obtain target-aware features that are robust to a multi-category scenario. Furthermore, we introduce a background token to effectively associate the target region of query with exemplars and decouple its background region from them. Our extensive experiments demonstrate that our model reaches a new state-of-the-art performance on the two challenging benchmarks, FSCD-LVIS and FSC-147, with a remarkably reduced degree of the target confusion problem.
@@ 1 Introduction
Object counting has achieved remarkable advances along with deep learning networks. However, most existing object counting methods are designed for a limited number of categories such as human macrowd or car car2. In fact, those methods highly rely on a large amount of labeled data and cannot handle unseen categories beyond training data. In this regard, few-shot object counting Lu et al. (2019) has been proposed to count arbitrary class objects in a query image based on the given exemplar images.
A mainstream of few-shot object counting is the extract-and-match approach Lu et al. (2019); Yang et al. (2021); Ranjan et al. (2021); Shi et al. (2022); Gong et al. (2022); You et al. (2023); Lin et al. (2022); Liu et al. (2022); Djukić et al. (2023); Gao and Huang (2024). Generally, this pipeline consists of three key components: 1) feature extractor, 2) relation learner, and 3) decoder. Firstly, they compute query and exemplar features using the feature extractor, then construct the correlation volume through the relation learner. Afterward, they estimate the number of instances in the query image by transferring the correlation volume to the decoder.
@@
Although previous studies You et al. (2023); Djukić et al. (2023) have achieved impressive performance, they exhibit a target confusion issue, failing to accurately identify only the target class when multiple classes of objects coexist in the query image, as shown in Figure 1. Existing methods have overlooked this problem, which is directly connected to the purpose of few-shot object counting, as benchmark datasets such as FSC-147 Ranjan et al. (2021) primarily consist of single-class scenes. The main reason for the target confusion is that the query features are computed without any explicit guidance of the target class. Consequently, the query features tend to focus on objectness rather than target class-specific features, hindering the differentiation between target and non-target object features.
To address this, we propose a novel framework, Mutually-Aware FEAture Learning (MAFEA), which enables the early consideration of mutual relations between query and exemplar features to produce the target class-specific features. Specifically, MAFEA employs cross-attention to capture bi-directional co-relations between query and exemplar features, along with self-attention to reflect internal relationships. With the cross-attention operating in a unified embedding space, the model can identify the difference between the target and non-target object features based on the exemplar features. However, in the cross-attention, query background features including non-target object features are inherently expressed by exemplar features since there are no other features except exemplar features. This might blur the distinction between the target and background features. To prevent this, we introduce a background token, which is incorporated alongside exemplar features in both self- and cross-attentions. This token is trained by our newly proposed Target-Background Discriminative (TBD) loss to effectively represent background features, including non-target object features. Consequently, MAFEA can capture the mutual relations beginning from the feature extractor and recognize the target objects clearly in the multi-category scenario as shown in Figure 1.
@@ 2.1 Class-Specific Object Counting
Class-specific object counting aims to count objects of a specific class, such as people macrowd, animals Arteta et al. (2016), and cars car2, in the images. For this purpose, traditional methods Leibe et al. (2005); Wang and Wang (2011); Stewart et al. (2016) solve the problem with detection-based approaches. Since most datasets provide only point-level supervision, most detection-based methods generate the pseudo-bounding boxes from point-level ground truth and update them in the training phase. However, they often struggle with scale variation and occlusion. To alleviate this problem, regression-based approaches Yan et al. (2019); Zhang et al. (2016); Wang et al. (2022, 2021) have emerged as popular alternatives in object counting, treating the task as dense regression to predict the object density map. This approach adeptly tackles overlap problems and achieves good performance. However, both of these approaches cannot handle object classes that are not present in the training phase.
2.2 Few-Shot Object Counting
Few-shot object counting aims to count arbitrary categories in the query image with just a few exemplar images. Pioneering methods Lu et al. (2019); Yang et al. (2021); Ranjan et al. (2021); Gong et al. (2022); Gao and Huang (2024) extract query and exemplar features, and leverage exemplar features as a fixed kernel to produce a correlation volume with query features. BMNet+ Shi et al. (2022) addresses a limitation in kernel-based matching mechanisms, which lack flexible channel interactions, by introducing a dynamic similarity metric capturing key exemplar patterns. SAFECount You et al. (2023) suggests a similarity-aware feature enhancement block that combines the advantages of both features and correlation volume. Recently, SPDCN Lin et al. (2022) and LOCA Djukić et al. (2023) integrate the shape and appearance properties of exemplars to reflect diverse object scales.
These methods have achieved impressive performances, but they do not account for the target confusion issue sufficiently. To alleviate this, we introduce a novel framework, Mutual-Aware Feature Learning, which computes the query and exemplar features dependently on each other throughout the process of feature extraction.
2.3 Vision Transformer
Motivated by the great success in the field of natural language processing, extensive studies have been conducted to employ self-attention for vision tasks such as image classification  Ramachandran et al. (2019); Dosovitskiy et al. (2021), object detection  Carion et al. (2020); Meng et al. (2021), and semantic segmentation  Yin et al. (2022); Caron et al. (2021). Also, the transformer-based encoder has advantages in class-specific object counting Tran et al. (2022); Liang et al. (2022). Recently, for few-shot object counting, CounTR  Liu et al. (2022) utilizes a transformer-based architecture to capture self-similarity prior explicitly and shows a powerful performance.
Unlike CounTR, which exploits separate feature extractors for query and exemplar images, MAFEA utilizes a mutually-aware feature extractor, computing the query and exemplar features in a unified embedding space. Moreover, MAFEA clearly distinguishes target object features from background features by introducing the background token.
@@ 3 Method
Figure 2: Comparison between Extract-and-Match methods and our proposed MAFEA. (a) Existing methods extract query and exemplar features without any explicit feedback to each other. (b) On the other hand, MAFEA produces the query and exemplar features based on their mutual relation from an early stage of the feature extractor. By integrating self-relations and bi-directional co-relations, MAFEA produces highly target-aware features. Moreover, the learnable background token is fed into the self- and co-relations with the exemplar features to represent the background regions of the query image.
@@ 3.2 Mutual Relation Modeling
$&s^{Q}=\text{MHA}(Q^{Q},K^{Q},V^{Q}),\\ &s^{E}=\text{MHA}(Q^{E},K^{E},V^{E}),\end{aligned}$ (1)
@@
$&c^{E\rightarrow{Q}}=\text{MHA}(Q^{Q},K^{E},V^{E}),\\ &c^{Q\rightarrow{E}}=\text{MHA}(Q^{E},K^{Q},V^{Q}).\end{aligned}$ (2)
@@
[…]
where $z^{Q}_{l}$ and $z^{E}_{l}$ are the input sequences of the $l$-th encoder layer. This modeling enables the encoder to adapt features not only based on their inherent self-relations but also their interrelated correlations.
@@ 3.3 Background Token
When computing $c^{E\rightarrow{Q}}$ in Eq. LABEL:eq:co-relation, MAFEA utilizes only exemplar features to produce keys and values. Although the attention mechanism intrinsically mitigates improper co-relations, the background features, including non-target object features, might be represented by the exemplar features. In this case, it obscures the difference between the target object and background features, thus, it confuses in precisely identifying the actual target in the query features. In this regard, we introduce the background token which is designed to learn the general features of the background region. The background token $z^{B}\in\mathbb{R}^{1\times C}$ is concatenated with the exemplar features and then fed into the self-relation and co-relation modeling as follows:
$&s^{[E;B]}=\text{MHA}([Q^{E};Q^{B}],[K^{E};K^{B}],[V^{E};V^{B}]),\\ &c^{[E;B]\rightarrow{Q}}=\text{MHA}(Q^{Q},[K^{E};K^{B}],[V^{E};V^{B}]),\\ &c^{Q\rightarrow{[E;B]}}=\text{MHA}([Q^{E};Q^{B}],K^{Q},V^{Q}),\end{aligned}$ (4)
@@ 3.4 Target-Background Discriminative Loss
Although the background token is designed to handle the background features of the query, it is not guaranteed without an explicit objective. In this regard, we define a target-background discriminative (TBD) loss which encourages the background token to align with background features. We first compute alignment score $AS_{i}$, which represents the degree of alignment between $i$-th query feature and the background token, as follows:
@@
where $Q^{Q}_{i}$ is $i$-th $Q^{Q}$, and $K^{{B}}_{j}$ and $[K^{E};K^{B}]_{j}$ denote $j$-th $K^{{B}}$ and $[K^{E};K^{B}]$, respectively. Then, to align the background token only with background features, we divide the query features into positive set $P$, comprising features that spatially contain one or more ground-truth (GT) points, and negative set $N$ which consists of features not including any GT points. As a result, we define TBD for the $i$-th query feature, as follows:
@@ 3.5 Training Loss
Once obtaining the target-aware features, $z^{Q}$ and $z^{E}$, we produce the correlation volume using the relation learner and convert it to the density map with the decoder. Following Djukić et al. (2023), the relation learner performs iterative adaptation to produce intermediate correlation volumes, subsequently processed by auxiliary decoder blocks as follows:
@@
$\mathcal{L}^{\text{count}}=\frac{1}{M}\left||y-\hat{y}|\right|_{2}^{2},$ (8)
@@
$\mathcal{L}^{\text{aux}}=\frac{1}{M}\sum_{k=1}^{K-1}\left||y_{k}-\hat{y}|\right|_{2}^{2},$ (9)
where $y_{k}$ is the intermediate density map of the $k$-th decoder and $K-1$ is the number of intermediate density maps. The full objectives are defined as follows:
@@ 4 Experiments
In this section, we first describe the experimental settings. We then compare it with the current state-of-the-art methods. Finally, we provide in-depth analyses of the results through the various ablation studies.
@@ 4.1 Implementation Details
Our framework comprises a ViT encoder, relation learner, and CNN decoder. The patch size is set to $16\times 16$, and both the kernel and stride of the projection head are set to $16\times 16$ corresponding to the patch size. The ViT encoder comprises 12 transformer encoder blocks where the hidden dimension of each block is 768, and the multi-head attention of each block consists of 12 heads. The relation learner, inspired by LOCA Djukić et al. (2023), incorporates Object Prototype Extraction (OPE) modules and Prototype Matching (PM). OPE integrates object shapes and appearance properties using a three-layered iterative adaption module. Each layer includes two multi-head attention blocks and a feed-forward network. Instead of ROI pooled features, we utilize exemplar features as appearance properties. PM involves the depth-wise correlation and max operation. Further details are provided in the appendix. The CNN decoder comprises 4 convolutions and bilinear upsampling layers to regress a 2D density map. The ViT encoder is initialized with a self-supervised pre-trained weight from MAEHe et al. (2022). On the other hand, the parameters of the relation learner and CNN decoder are randomly initialized.
@@
We apply the same pre-processing as LOCA. As the inputs, the query image is resized to $512\times 512$ and exemplar images are scaled to $48\times 48$ based on provided box annotations. We set the weights $\lambda_{1}$ and $\lambda_{2}$ for auxiliary and TBD loss in Eq. 10 to $0.3$ and $0.05$, respectively. AdamW optimizer is employed with a batch size of 8. The initial learning rate is $1e^{-4}$ and is halved every 40 epochs. Training is performed on a single RTX3090 GPU for 100 epochs.
@@
We experiment on three benchmark datasets: FSCD-LVIS Nguyen et al. (2022), FSC-147 Ranjan et al. (2021), and CARPK Hsieh et al. (2017). FSCD-LVIS is designed for few-shot object counting and detection in complex scenes with multiple class objects. It contains $6195$ images across $372$ classes, split into $3953$ train and $2242$ test images. FSC-147, a few-shot object counting dataset, consists of simpler scenes where most images contain only a target class. It includes $6135$ images of $147$ classes, divided into $3659$ train, $1286$ validation, and $1190$ test images. Both datasets include three randomly selected exemplars per image to depict the target objects. Note that, there is no shared object class between the sets. Furthermore, we validate our models’ generalization capability on the test set of CARPK, a dataset tailored for counting cars, comprising $459$ drone-captured images.
@@
Although the FSC-147 contains a large number of objects in each image, the scene of each image is mostly composed of single-class objects. Due to this inherent characteristic, the evaluation of the FSC-147 might not accurately assess the ability of the model to identify the target class within an image containing diverse object categories. For a quantitative assessment of whether the model suffers from the target confusion problem, we construct a multi-class subset of the FSC-147 (FSC-147-Multi). We selectively remove images where objects from other classes amount to less than 20% of the target class objects, to exclude single-object predominant images in the multi-class experiments. The indices of the images that make up the FSC-147-Multi can be found in Table 1, and experimental results are detailed in Sec. 4.3.
@@
Table 2: Evaluation on a multi-class scenario with 3-shot exemplars. The results on FSCD-LVIS are reproduced using the official implementation, employing identical query image sizes of $512$. The results on the multi-class subset of FSC-147 (FSC-147-Multi) are evaluated using the official pre-trained weights. The subset consists of 31 (over 1286) in the validation set and 12 (over 1190) images in the test set, totaling 43 images. ‘-’ means that the score is not available. $\dagger$’ means the mean and standard deviation of the results over five runs with $95\%$ confidence intervals.
Method FSCD-LVIS FSC-147-Multi
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Counting-DETR Nguyen et al. (2022) $23.50$ $35.89$
BMNet+ Shi et al. (2022) $17.04$ $27.71$ $11.22$ $16.99$
SPDCN Lin et al. (2022) $14.36$ $26.31$ $9.27$ $13.54$
CounTR Liu et al. (2022) $14.14$ $26.01$ $16.12$ $30.70$
SAFECount You et al. (2023) $14.01$ $24.04$ $8.63$ $12.18$
LOCA Djukić et al. (2023) $14.38$ $24.28$ $9.97$ $16.83$
MAFEA (Ours) 12.47 22.94 6.36 9.18
LOCA Djukić et al. (2023) $14.51\pm 0.12$ $24.70\pm 0.88$ $10.41\pm 0.69$ $17.29\pm 1.63$
MAFEA (Ours) $\textbf{12.51}\pm 0.09$ $\textbf{23.11}\pm 0.09$ $\textbf{5.52}\pm 0.66$ $\textbf{7.93}\pm 1.06$
Table 3: Evaluation on a single-class scenario with 3-shot exemplars in FSC-147 dataset. $\dagger$’ indicates the mean and standard deviation of the results over five runs with $95\%$ confidence intervals.
Method Val Test
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
GMN Lu et al. (2019) $29.66$ $89.81$ $26.52$ $124.57$
FamNet Ranjan et al. (2021) $24.32$ $70.94$ $22.56$ $101.41$
CFOCNet Yang et al. (2021) $21.19$ $61.41$ $22.10$ $112.71$
RCAC Gong et al. (2022) $20.54$ $60.78$ $20.21$ $81.86$
Counting-DETR Nguyen et al. (2022) - - $16.79$ $123.56$
BMNet+ Shi et al. (2022) $15.74$ $58.53$ $14.62$ $91.83$
SPDCN Lin et al. (2022) $14.59$ $49.97$ $13.51$ $96.80$
CounTR Liu et al. (2022) $13.13$ $49.83$ $11.95$ $91.23$
SAFECount You et al. (2023) $15.28$ $47.20$ $14.32$ $85.54$
LOCA Djukić et al. (2023) $10.24$ $32.56$ $10.79$ $56.97$
CSTrans Gao and Huang (2024) $18.10$ $58.45$ $16.38$ $93.51$
MAFEA (Ours) 8.92 32.45 9.84 56.68
LOCA Djukić et al. (2023) $10.57\pm 0.19$ $34.30\pm 1.79$ $11.39\pm 0.32$ $72.82\pm 7.16$
MAFEA (Ours) $\textbf{9.24}\pm 0.22$ $\textbf{33.88}\pm 0.74$ $\textbf{10.23}\pm 0.40$ $\textbf{64.96}\pm 5.65$

### v2

@@ Abstract
Few-shot object counting has garnered significant attention for its practicality as it aims to count target objects in a query image based on given exemplars without additional training. However, the prevailing extract-and-match approach has a shortcoming: query and exemplar features lack interaction during feature extraction since they are extracted independently and later correlated based on similarity. This can lead to insufficient target awareness and confusion in identifying the actual target when multiple class objects coexist. To address this, we propose a novel framework, Mutually-Aware FEAture learning (MAFEA), which encodes query and exemplar features with mutual awareness from the outset. By encouraging interaction throughout the pipeline, we obtain target-aware features robust to a multi-category scenario. Furthermore, we introduce background token to effectively associate the query’s target region with exemplars and decouple its background region. Our extensive experiments demonstrate that our model achieves state-of-the-art performance on FSCD-LVIS and FSC-147 benchmarks with remarkably reduced target confusion.
@@ 1 Introduction
Object counting has achieved remarkable advances along with deep learning networks. However, most existing object counting methods are designed for a limited number of categories, such as human macrowd or car car2. Those methods highly rely on a large amount of labeled data and cannot handle unseen categories beyond training data. In this regard, few-shot object counting GMN has been proposed to count arbitrary class objects in a query image based on the given exemplar images.
Recent works CounTR; LOCA have explored both the few-shot and zero-shot settings. The few-shot setting counts target objects based on exemplar images, while the zero-shot setting counts the most frequent class by identifying repetitive patterns without explicit guidance. While both few-shot and zero-shot settings offer valuable approaches, few-shot setting plays a critical role in accurately defining the object unit in complex real-world applications. This is because recognizing repetitive patterns alone may not provide enough clarity regarding the unit of the object to count, whereas visual exemplars inherently contain information about the object unit.
A mainstream of few-shot object counting is the extract-and-match approach GMN; CFOCNet; FamNet; BMNet+; RCAC; SAFECount; SPDCN; CounTR; LOCA; cstrans. Generally, this pipeline consists of three key components: 1) feature extractor, 2) relation learner, and 3) decoder. Firstly, they compute query and exemplar features using the feature extractor, then construct the correlation volume through the relation learner. Afterward, they estimate the number of instances in the query image by transferring the correlation volume to the decoder.
@@
Although previous studies SAFECount; LOCA have achieved impressive performance, they exhibit a target confusion issue, failing to accurately identify only the target class when multiple classes of objects coexist in the query image, as shown in Figure 1. Existing methods have overlooked this problem, directly connected to the purpose of few-shot object counting, as benchmark datasets such as FSC-147 FamNet primarily consist of single-class scenes. The main reason for the target confusion is that the query features are computed without any explicit guidance of the target class. Consequently, the query features tend to focus on objectness rather than target class-specific features, hindering the differentiation between target and non-target object features.
To address this, we propose a novel framework, Mutually-Aware FEAture Learning (MAFEA), which enables the early consideration of mutual relations between query and exemplar features to produce the target class-specific features. Specifically, MAFEA employs cross-attention to capture bi-directional co-relations between query and exemplar features, along with self-attention to reflect internal relationships. With the cross-attention operating in a unified embedding space, the model can identify the difference between the target and non-target object features based on the exemplar features. However, in the cross-attention, query background features, including non-target object features, are inherently expressed by exemplar features since there are no other features except exemplar features. This might blur the distinction between the target and background features. To prevent this, we introduce a background token, incorporated alongside exemplar features in self- and cross-attentions. Our newly proposed Target-Background Discriminative (TBD) loss trains this token to effectively represent background features, including non-target object features. Consequently, MAFEA can capture the mutual relations beginning from the feature extractor and recognize the target objects clearly in the multi-category scenario, as shown in Figure 1.
@@ 2.1 Class-Specific Object Counting
Class-specific object counting aims to count objects of a specific class, such as people macrowd, animals animal, and cars car2, in the images. For this purpose, traditional methods class-specific; class-specific2; class-specific3 solve the problem with detection-based approaches. Since most datasets provide only point-level supervision, most detection-based methods generate the pseudo-bounding boxes from point-level ground truth and update them in the training phase. However, they often struggle with scale variation and occlusion. To alleviate this problem, regression-based approaches class-specific-regression; class-specific-regression2; STNet; Interlayer have emerged as popular alternatives in object counting, treating the task as dense regression to predict the object density map. This approach adeptly tackles overlap problems and achieves good performance. However, neither of these approaches can handle object classes not present in the training phase.
2.2 Few-Shot and Zero-Shot Object Counting
Few-shot object counting aims to count objects from arbitrary categories in a query image using only a few exemplar images. Previous methods GMN; CFOCNet; FamNet; RCAC; cstrans; BMNet+; fewshotseg; SAFECount; SPDCN; CounTR; LOCA extract query and exemplar features independently, and match these features to infer a density map. Recent methods such as CounTR CounTR and LOCA LOCA have evaluated zero-shot ability by replacing exemplar images with learnable tokens, extending beyond the few-shot setting. GCNet GCNet, on the other hand, proposed an exemplar-free approach and expanded it with weak supervision at the count level.
In parallel, zero-shot object counting has emerged, which enables counting based on text input instead of exemplar images. ZSCNet ZSCNet was the first to introduce a method that identifies objects of interest using class names. More recently, SAM-Free SAM-Free and PseCo PseCo have attempted to leverage the powerful generalization abilities of foundation models like SAM SAM and CLIP CLIP. However, since SAM is trained for segmentation and CLIP for classification, these approaches struggle in heavy occlusion, small objects, and dense scenes due to task discrepancies.
It has been assumed that current few-shot and zero-shot counting methods can function in multi-class scenarios. However, this has yet to be proved, as the benchmark dataset FamNet contains only one class per image. We experimentally confirmed that existing methods struggle with target confusion in multi-class scenes, whether in few- or zero-shot settings. We conjecture that target confusion arises because existing methods encode query and exemplar features independently, without any explicit interaction during feature extraction. This lack of interaction causes insufficient target awareness, making it difficult for the model to distinguish between target and non-target objects in multi-class scenes. To settle this problem, we designed the model to distinguish between target and non-target objects by producing target-aware features through early interaction in the feature extraction process. To the best of our knowledge, this is the first attempt to highlight and address the issue of target confusion.
@@ 3 Method
Figure 2: Comparison between Extract-and-Match methods and our proposed MAFEA. (a) Existing methods extract query and exemplar features without any explicit feedback to each other. (b) On the other hand, MAFEA produces the query and exemplar features based on their mutual relation from an early stage of the feature extractor. By integrating self-relations and bi-directional co-relations, MAFEA produces highly target-aware features. Moreover, the learnable background token is fed into the self- and co-relations with the exemplar features to represent the background regions of the query image. (c) Self- and Cross-Relations are implemented by self- and cross-attention mechanisms.
@@ 3.2 Mutual Relation Modeling
$\begin{aligned} &s^{Q}=\text{MHA}(Q^{Q},K^{Q},V^{Q}),\\ &s^{E}=\text{MHA}(Q^{E},K^{E},V^{E}),\end{aligned}$ (1)
@@
$\begin{aligned} &c^{E\rightarrow{Q}}=\text{MHA}(Q^{Q},K^{E},V^{E}),\\ &c^{Q\rightarrow{E}}=\text{MHA}(Q^{E},K^{Q},V^{Q}).\end{aligned}$ (2)
@@
$\begin{aligned} &z^{Q}_{l+1}=z^{Q}_{l}+s^{Q}_{l}+c^{E\rightarrow{Q}}_{l},\quad l=1,2,\dots,L-1,\\ &z^{E}_{l+1}=z^{E}_{l}+s^{E}_{l}+c^{Q\rightarrow{E}}_{l},\quad l=1,2,\dots,L-1,\end{aligned}$ (3)
where $z^{Q}_{l}$ and $z^{E}_{l}$ are the input sequences of the $l$-th encoder layer. This modeling enables the encoder to adapt features based on their inherent self-relations and their interrelated correlations.
@@ 3.3 Background Token
When computing $c^{E\rightarrow{Q}}$ in Eq. LABEL:eq:co-relation, MAFEA utilizes only exemplar features to produce keys and values. Although the attention mechanism intrinsically mitigates improper co-relations, the exemplar features might represent the background features, including non-target object features. In this case, it obscures the difference between the target object and background features; thus, it confuses the precise identification of the target in the query features. In this regard, we introduce the background token, designed to learn the general features of the background region. The background token $z^{B}\in\mathbb{R}^{1\times C}$ is concatenated with the exemplar features and then fed into the self-relation and co-relation modeling as follows:
$\begin{aligned} &s^{[E;B]}=\text{MHA}([Q^{E};Q^{B}],[K^{E};K^{B}],[V^{E};V^{B}]),\\ &c^{[E;B]\rightarrow{Q}}=\text{MHA}(Q^{Q},[K^{E};K^{B}],[V^{E};V^{B}]),\\ &c^{Q\rightarrow{[E;B]}}=\text{MHA}([Q^{E};Q^{B}],K^{Q},V^{Q}),\end{aligned}$ (4)
@@ 3.4 Target-Background Discriminative Loss
Although the background token is designed to handle the background features of the query, it is not guaranteed without an explicit objective. In this regard, we define a target-background discriminative (TBD) loss, which encourages the background token to align with background features. We first compute alignment score $AS_{i}$, which represents the degree of alignment between $i$-th query feature and the background token, as follows:
@@
where $Q^{Q}_{i}$ is $i$-th $Q^{Q}$, and $K^{{B}}_{j}$ and $K^{E};K^{B}]_{j}$ denote $j$-th $K^{{B}}$ and $K^{E};K^{B}]$, respectively. Then, to align the background token only with background features, we divide the query features into positive set $P$, comprising features that spatially contain one or more ground-truth (GT) points, and negative set $N$ which consists of features not including any GT points. As a result, we define TBD for the $i$-th query feature, as follows:
@@ 3.5 Training Loss
Once obtaining the target-aware features, $z^{Q}$ and $z^{E}$, we produce the correlation volume using the relation learner and convert it to the density map with the decoder. Following LOCA, the relation learner performs iterative adaptation to produce intermediate correlation volumes, subsequently processed by auxiliary decoder blocks as follows:
@@
[…]
@@
[…]
where $y_{k}$ is the intermediate density map of the $k$-th decoder and K1K-1 is the number of intermediate density maps. The full objectives are defined as follows:
@@ 4.1 Implementation Details
Our framework comprises a ViT encoder, relation learner, and CNN decoder. The patch size is set to $16\times 16$, and both the kernel and stride of the projection head are set to $16\times 16$ corresponding to the patch size. The ViT encoder comprises 12 transformer encoder blocks where the hidden dimension of each block is 768, and the multi-head attention of each block consists of 12 heads. The relation learner, inspired by LOCA LOCA, incorporates Object Prototype Extraction (OPE) modules and Prototype Matching (PM). OPE integrates object shapes and appearance properties using a three-layered iterative adaption module. Each layer includes two multi-head attention blocks and a feed-forward network. Instead of ROI pooled features, we utilize exemplar features as appearance properties. PM involves the depth-wise correlation and max operation. Further details are provided in the appendix. The CNN decoder comprises 4 convolutions and bilinear upsampling layers to regress a 2D density map. The ViT encoder is initialized with a self-supervised pre-trained weight from MAEMAE. On the other hand, the parameters of the relation learner and CNN decoder are randomly initialized.
@@
We apply the same pre-processing and augmentation strategies as LOCA. For pre-processing, the query image is resized to $512\times 512$ and exemplar images are scaled to $48\times 48$ based on provided box annotations. Tiling augmentation is then applied, where the image is split into multiple tiles, and random horizontal flipping and color jitter are applied to each tile. In the zero-shot setting, where no exemplar is provided, we follow the same approach as LOCA by replacing exemplar images with a learnable token for training and evaluation. The weights $\lambda_{1}$ and $\lambda_{2}$ for auxiliary and TBD loss in Eq. 10 are set to $0.3$ and $0.05$, respectively. AdamW optimizer is employed with a batch size of 8, an initial learning rate of $1e^{-4}$, which is halved every 40 epochs, for a total of 100 epochs. All experiments are conducted on a single RTX3090 GPU, with a training time of approximately 10 hours and an inference time of 0.02 seconds per sample.
@@
We experiment on three benchmark datasets: FSCD-LVIS Counting-DETR, FSC-147 FamNet, and CARPK CARPK. FSCD-LVIS is designed for few-shot object counting and detection in complex scenes with multiple class objects. It contains $6195$ images across $372$ classes, split into $3953$ train and $2242$ test images. FSC-147, a few-shot object counting dataset, consists of simpler scenes where most images contain only a target class. It includes $6135$ images of $147$ classes, divided into $3659$ train, $1286$ validation, and $1190$ test images. Both datasets include three randomly selected exemplars per image to depict the target objects. Note that, there is no shared object class between the sets. Furthermore, we validate our models’ generalization capability on the test set of CARPK, a dataset tailored for counting cars, comprising $459$ drone-captured images.
@@
Table 2: Evaluation on multi-class datasets in zero- and few-shot settings. The results on FSCD-LVIS are reproduced using the official implementation, except for Counting-DETR, which uses its official performance. The results on the multi-class subset of FSC-147 (FSC-147-Multi) are evaluated using the official pre-trained weights. The subset consists of 31 (over 1286) in the validation set and 12 (over 1190) images in the test set, totaling 43 images. ‘-’ means that the score is not available and $\times$’ means that no exemplar is provided.
Method Exemplar FSCD-LVIS FSC-147-Multi
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
SAM-Free SAM-Free class name - - $17.79$ $26.74$
PseCo PseCo class name - - $14.74$ $23.23$
CounTR CounTR $\times$ - - $14.17$ $28.80$
LOCA LOCA $\times$ - - $12.91$ $24.39$
MAFEA (Ours) $\times$ $14.72$ $27.42$ $12.12$ $26.51$
Counting-DETR Counting-DETR box $23.50$ $35.89$ - -
BMNet+ BMNet+ box $17.04$ $27.71$ $11.22$ $16.99$
SPDCN SPDCN box $14.36$ $26.31$ $9.27$ $13.54$
CounTR CounTR box $14.14$ $26.01$ $16.12$ $30.70$
SAFECount SAFECount box $14.01$ $24.04$ $8.63$ $12.18$
LOCA LOCA box $14.38$ $24.28$ $9.97$ $16.83$
SAM-Free SAM-Free box - - $13.12$ $18.20$
PseCo PseCo box - - $19.81$ $38.64$
MAFEA (Ours) box 12.47 22.94 6.36 9.18
Table 3: Evaluation on the FSC-147 dataset in few- and zero-shot settings.
Method Exemplar Val Test
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
ZSCNet ZSCNet class name $26.93$ $88.63$ $22.09$ $115.17$
SAM-Free SAM-Free class name - - $24.79$ $137.15$
PseCo PseCo class name $23.90$ $100.33$ $16.58$ $129.77$
CounTR CounTR $\times$ $17.40$ $70.33$ $14.12$ $108.01$
LOCA LOCA $\times$ $17.43$ $54.96$ $16.22$ $103.96$
GCNet GCNet $\times$ $19.50$ $63.13$ $17.83$ $102.89$
MAFEA (Ours) $\times$ $14.54$ $60.44$ $13.23$ $105.99$
GMN GMN box $29.66$ $89.81$ $26.52$ $124.57$
FamNet FamNet box $24.32$ $70.94$ $22.56$ $101.41$
CFOCNet CFOCNet box $21.19$ $61.41$ $22.10$ $112.71$
RCAC RCAC box $20.54$ $60.78$ $20.21$ $81.86$
Counting-DETR Counting-DETR box - - $16.79$ $123.56$
BMNet+ BMNet+ box $15.74$ $58.53$ $14.62$ $91.83$
SPDCN SPDCN box $14.59$ $49.97$ $13.51$ $96.80$
CounTR CounTR box $13.13$ $49.83$ $11.95$ $91.23$
SAFECount SAFECount box $15.28$ $47.20$ $14.32$ $85.54$
LOCA LOCA box $10.24$ $32.56$ $10.79$ $56.97$
SAM-Free SAM-Free box - - $19.95$ $132.16$
GCNet GCNet box $19.61$ $66.22$ $17.86$ $106.98$
CSTrans cstrans box $18.10$ $58.45$ $16.38$ $93.51$
PseCo PseCo box $15.31$ $68.34$ $13.05$ $112.86$
MAFEA (Ours) box 8.92 32.45 9.84 56.68
@@
Although the FSC-147 contains many objects in each image, the scene of each image is mainly composed of single-class objects. Due to this inherent characteristic, the evaluation of the FSC-147 might not accurately assess the ability of the model to identify the target class within an image containing diverse object categories. For a quantitative assessment of whether the model suffers from the target confusion problem, we construct a multi-class subset of the FSC-147 (FSC-147-Multi). We selectively remove images where objects from other classes amount to less than 20% of the target class objects, to exclude single-object predominant images in the multi-class experiments. The indices of the images that make up the FSC-147-Multi can be found in Table 1, and experimental results are detailed in Sec. 4.3.